# Claim Verification in the Age of Large Language Models: A Survey

### v1

@@ Abstract
The large and ever-increasing amount of data available on the Internet coupled with the laborious task of manual claim and fact verification has sparked the interest in the development of automated claim verification systems. [^1] We use the terms claim verification and fact verification interchangeably given the overlap between the two concepts. Several deep learning and transformer-based models have been proposed for this task over the years. With the introduction of Large Language Models (LLMs) and their superior performance in several NLP tasks, we have seen a surge of LLM-based approaches to claim verification along with the use of novel methods such as Retrieval Augmented Generation (RAG). In this survey, we present a comprehensive account of recent claim verification frameworks using LLMs. We describe the different components of the claim verification pipeline used in these frameworks in detail including common approaches to retrieval, prompting, and fine-tuning. Finally, we describe publicly available English datasets created for this task.
Introduction
@@
Figure 2: The archetypal claim verification pipeline and its main components.
False information is widely present on social media and on the Web motivating the development of automated fact verification systems (Guo, Schlichtkrull, and Vlachos 2022). The introduction of LLMs has provided malicious actors with sophisticated ways of creating and disseminating false information. Recent election cycles saw a large number of claims spread across both social media and news platforms alike (Dmonte et al. 2024). Similarly, during the COVID-19 pandemic, many claims were spread across social media platforms. Many of these claims were factually inaccurate which led to the spread of misinformation (Zhou et al. 2023).
Fact-checking potentially false claims is an essential moderation task to reduce the spread of misinformation. Several organizations such as FactCheck, PolitiFact, NewsGuard, and Full Fact perform manual fact-checking to verify claims in different domains. However, this has been regarded as a laborious task that requires domain expertise (Adair et al. 2017; Hanselowski 2020) and it has become less and less feasible due to the sheer volume of misinformation that can be generated by humans and AI models. Automated fact-checking has become an increasingly popular approach to verify the veracity of claims in a given text. [^2] We acknowledge that claim verification models can also be applied to other modalities of data (e.g., images). In this survey, we address models applied to text only. There are several steps involved in the fact-verification pipelines, but the three main components are claim detection, evidence retrieval, and veracity prediction. Models used in these steps have followed the general methodological developments of the field and we have thus observed an increase in the use of LLMs for claim verification (Zhang and Gao 2023; Wang and Shu 2023; Quelle and Bovet 2024). Figure 1 shows an example claim veracity scenario and compares the architectural differences between NLP-based (traditional) and LLM-based claim verification systems. Compared to traditional systems which use NLP-based models, fact verification are less prone to error propagation, and provide explainable/interpretable generations/justification when verifying claims.
@@
Despite this, LLMs are pre-trained on very large collections of texts and are prone to hallucinations, often generating texts containing incorrect information. Furthermore, such models can be used to generate misinformation at scale (Chen and Shu 2023; Zhou et al. 2023; Dmonte et al. 2024) and therefore they can be exploited by malicious actors to generate and spread wrong and factually incorrect information at an unprecedented rate (Pan et al. 2023c). Furthermore, using these pre-trained models for fact-verification may generate incorrect veracity labels, as the models may also rely on obsolete information to assess the veracity of a claim. Approaches like RAG (Gao et al. 2023), which aid the models in their decision-making abilities are used in this task for the model to get access to the most recent information during the fact-verification task.
@@
A few general automated claim verification surveys have been published over the years (Zeng, Abumansour, and Zubiaga 2021; Bekoulis, Papagiannopoulou, and Deligiannis 2021; Guo, Schlichtkrull, and Vlachos 2022) as well as a couple of surveys focusing on particular aspects of the task such as explainability (Vallayil et al. 2023) and applications to specialized domains such as scientific texts (Vladika and Matthes 2023). However, all past related surveys, lack consideration for LLM-based approaches or focus on specific sub-tasks of the pipeline (Panchendrarajan and Zubiaga 2024). In this paper, we fill this important gap by surveying LLM-based frameworks proposed in recent years. To the best of our knowledge, this is the first survey to explore claim verification with LLMs. We expect it to be a valuable resource to researchers in the field opening exciting new avenues for future research.
Search Criteria
@@
We search various well-known repositories [^3] ACM: portal.acm.org. IEEE Xplore: ieeexplore.ieee.org. Scopus: scopus.com. ACL: aclanthology.org. Web of Science: isiknowledge.com. Springer: link.springer.com. ArXiv: arxiv.org. CEUR: https://ceur-ws.org/. of scientific articles to collect the papers that serve as primary sources for this survey. We apply queries including LLMs, claim verification, fact verification and related keywords on these repositories. We focus primarily on the ACL Anthology, the ACM Digital Library, and IEEE Xplore, and proceedings of related conferences such as AAAI and IJCAI. We further search on Scopus, Springer Link, and Science@Direct, and ArXiv.
@@
We collected and reviewed over 100 papers based on our search terms and filter papers based on the following search criteria: (i) since the scope of our review is on text-based LLM claim verification systems, we omit papers related to multimodal approaches (e.g., text-to-images), our papers from other modalities in claim verification (e.g., images, graph-based systems). (ii) since our review focuses on LLM-based systems for claim veracity, we omit papers related to claim identification, claim detection, and/or detecting LLM-generated content. Based on our filtering criteria, we collected a total of 49 papers related to LLM-based approaches for veracity labeling. The papers on topic LLM-based veracity labeling have been published primarily in the ACL Anthology and ACM Digital Library, but also in other repositories such as the IEEE Xplore. We release this curated list to be publicly available.
Figure 3: Claim verification pipeline using large language models. Instead of traditional automatic claim verification, this pipeline creates a prompt from the retrieved evidence and the input claim as input to the LLM to generate a label, sentence evidence, and/or explanation of its response.
Claim Verification Pipeline
@@
Figure 2 shows an archetypal claim verification pipeline consisting of the following modules: claim detection, claim matching, claim check-worthiness, document/evidence retrieval, rationale/sentence selection, veracity label prediction, and explanation/justification generation. Many proposed systems make use of only some of these modules. The following subsections describe each of these modules in detail. Finally, as described in Panchendrarajan and Zubiaga (2024), performance evaluation in these sub-tasks is usually carried out using well-established automatic evaluation metrics used in text classification such as Precision, Recall, and F-score.
@@
Input texts may contain one or more statements but not all statements are claims. Given the input text, a claim detection module is designed to identify all the statements containing a claim. For example, the statement ’I loved the movie Oppenheimer.’ is an opinion, whereas the statement ’The COVID-19 pandemic started in Texas.’ contains a claim.
@@
Not all the identified claims are check-worthy. In the check-worthy claim identification sub-task, given an input claim, the task is to identify the claims that include real-world assertions and that may need to be verified (Hassan, Li, and Tremayne 2015; Nakov et al. 2021). This is a subjective task and it relies on the factors like popularity of the claim, public interest in determining the veracity of the claim, etc. For example, the claim ’The President met the State Governor to discuss the infrastructure deal.’ is less check-worthy than the claim ’Drinking salt water cures COVID.’.
@@
The identified check-worthy claims can be matched to the previously fact-checked claims. Given an input claim and a database of previously fact-checked claims, claim matching is used to determine if the input claim is previously fact-checked and exists in the database (Shaar et al. 2020; Nakov et al. 2021). This can help avoid the next steps in the pipeline and a veracity label can be predicted directly.
@@
If the input claim does not exist in the database of fact-checked claims, the claim needs to be verified. In the document or evidence retrieval sub-task, all the relevant documents related to the input claim are extracted, either from an external database or through an internet search/information retrieval(Chen et al. 2017). A threshold of how many documents are to be retrieved can be predetermined.
@@
All the information in the retrieved documents is not relevant to the claim. Hence in the rationale selection task, only the information or evidence most relevant to the claim are selected to be used to predict the veracity label.
@@
Once the rationale is selected, the claim along with the rationale and additional features if any, are given as input to a machine learning model. The task of the classifier is to predict a veracity label from among the following three labels; ’SUPPORTED’, ’REFUTED’, or ’NOT ENOUGH EVIDENCE’. In some datasets, the labels can also be ’TRUE’, or ’FALSE’, depending on the specific task.
@@
Recent works have focused on generating explanations for the veracity labels prediction. This specific task is focused on generating natural language justifications or explanations for the prediction considering the claim and evidence to generate these explanations.
LLM Approaches
@@
In this section, we describe the recent advancements that enable LLMs to be robust in fact verification scenarios. Figure 3 shows an example pipeline that encapsulates multiple component modules (i.e., Evidence Retrieval, Prompt Creation, Transfer Learning, and LLM Generation) for verifying claims. Different from traditional fact verification pipelines that select evidence set for verifying claims, LLM-based claim verification conditions generated text based on the concatenated input claim and retrieved evidence. The ability of large, pre-trained LMs augmented with retrieval enables them to perform well on knowledge-intensive tasks such as text generation.
Evidence Retrieval Strategies
@@
RAG models, which have been developed to combat the issue of hallucination in LLMs in knowledge-intensive tasks, have shown success in the scope of fact verification (Lewis et al. 2020; Izacard et al. 2023; Gao et al. 2023; Guan et al. 2023). Early work such as in Lewis et al. (2020) developed a framework that incorporates a retriever to retrieve evidence from an external database such as Wikipedia for conditionally generating veracity labels in fact verification. Authors in Izacard et al. (2023) demonstrate that RAGs perform well on the FEVER shared task (Thorne et al. 2018b)) in few-shot settings, showing approximately 5% improvement over large-scale LLMs such as Gopher (Rae et al. 2021) with significantly fewer parameters (i.e., 11B compared to 280B). Other works consider the optimization of either document ranking (Glass et al. 2022; Chen et al. 2022c; Hofstätter et al. 2023) or input claims (i.e., queries) (Hang, Yu, and Tan 2024) as a crucial step for improving evidence retrieval for veracity labeling. Authors in Hofstätter et al. (2023) use an autoregressive re-ranker to get the most relevant passages from the retriever. These are then passed to the generation model to generate the veracity label.
@@
Despite this, RAG models can often fail when encountering long or complex input claims, causing the model to incorrectly generate veracity labels or evidence sentences. Recent works have addressed this issue by segmenting long claims into smaller sub-claims and performing multiple rounds of retrieval (Khattab, Potts, and Zaharia 2021; Shao et al. 2023). Authors in Khattab, Potts, and Zaharia (2021) present a pipeline for multi-hop claim verification that uses an iterative retriever and neural methods for effective document retrieval and re-ranking.  Shao et al. (2023) show that using a re-ranker to distill knowledge to a retriever helps close the semantic gaps between a query and document passage when verifying claims using iterative RAGs. Other works address the issue of complex claims by using fine-grained retrieval techniques based on claim decomposition (Chen et al. 2023a; Zhang and Gao 2023; Pan et al. 2023b).Chen et al. (2023a) generate sub-questions based on a claim, which a document retriever uses to retrieve relevant documents. A fine-grained retriever retrieves top-k text spans as evidence based on a k-word window and BM25. Authors in Zhang and Gao (2023) decompose a claim into sub-claims and generate questions to verify the sub-claims. External knowledge sources are used to retrieve relevant information to verify the sub-claims and generate a final veracity label. Pan et al. (2023b) follow a programming paradigm, where the claim is broken down into subtasks and the final label is the aggregation of the execution of each subtask. The fact-verification subtask uses external knowledge to retrieve relevant evidence for a claim.
@@
Hang, Yu, and Tan (2024) retrieve evidence based on generated knowledge graphs of evidences. They generate a knowledge graph of user query or input claim and compare it to the database of knowledge graphs to retrieve the most relevant information for claim verification. Authors in Hu et al. (2023) propose a latent variable model that allows the retrieval of the most relevant evidence sentences from a document while removing the irrelevant sentences. This approach reduces the noisy data during the verification process. Stochastic-RAG, an approach proposed by Zamani and Bendersky (2024) uses a stochastic sampling without replacement process for evidence retrieval and selection. This approach overcomes the ranking and selection of the evidence hence optimizing the RAG model. Authors in Zhang et al. (2023) optimize the evidence retrieval process by using feedback from the claim verifier. The divergence between the evidence from a retrieved evidence set provided to the verifier and the gold standard evidence, acts as a feedback signal used to train the retriever. Xu et al. (2024b) propose Search-in-the-Chain, an approach where an LLM generates a reasoning chain and based on the answer to each node in the chain, retrieval can be used to correct an answer or provide additional knowledge. The approach improves the generation accuracy of the LLM.
Prompt Creation Strategies
@@
Text prompting has been shown to be an effective technique for improving the desired output of large-scale generative models. In the context of claim verification, several works investigate prompting strategies, using both manual and automated techniques, for building more robust claim verification systems Zhang and Gao (2023); Li et al. (2023c); Zeng and Gao (2023); Chen et al. (2023b). Authors in (Zhang and Gao 2023) develop a hierarchical prompting technique that enables LLMs to verify multiple sub-claims using a step-by-step approach. The work in Li et al. (2023c) demonstrates a self-sufficient claim verification through prompting instructions on multiple language models.
@@
ProToCo (Zeng and Gao 2023) demonstrates improved claim verification performance of LLMs by leveraging a consistency mechanism to construct variants of the original claim-evidence pair prompt based on three logical relations (i.e., confirmation, negation, uncertainty). Authors in Chen et al. (2023b) develop a unified retrieval framework that employs discrete, continuous, and hybrid prompt strategies for adjusting to various knowledge-intensive tasks such as claim verification. Other works such as the FactualityPrompts (Lee et al. 2022) framework test the output generations of LLMs given an input prompt and use an external database such as Wikipedia to calculate factuality and quality measures compared to the ground truth. Other works aim to improve the LLM’s reasoning abilities by appending the claims with evidence during prompting in the context of claim verification and text generation (Parvez 2024; Dougrez-Lewis et al. 2024).
Transfer Learning Strategies

### v2

@@ Abstract
The large and ever-increasing amount of data available on the Internet coupled with the laborious task of manual claim and fact verification has sparked the interest in the development of automated claim verification systems. [^1] We use the terms claim verification and fact verification interchangeably given the overlap between the two concepts. Several deep learning and transformer-based models have been proposed for this task over the years. With the introduction of Large Language Models (LLMs) and their superior performance in several NLP tasks, we have seen a surge of LLM-based approaches to claim verification along with the use of novel methods such as Retrieval Augmented Generation (RAG). In this survey, we present a comprehensive account of recent claim verification frameworks using LLMs. We describe the different components of the claim verification pipeline used in these frameworks in detail including common approaches to retrieval, prompting, and fine-tuning. Finally, we describe publicly available English datasets for this task.
1 Introduction
False information is widely present on social media and on the Web motivating the development of automated fact verification systems Guo et al. (2022). The introduction of LLMs has provided malicious actors with sophisticated ways of creating and disseminating false information. Recent election cycles has seen a large number of claims spread across social media and news platforms alike Dmonte et al. (2024). Similarly, during the COVID-19 pandemic, many factually inaccurate claims were spread Zhou et al. (2023).
Fact-checking potentially false claims is an essential moderation task to reduce the spread of misinformation. Organizations such as FactCheck, PolitiFact, NewsGuard, and Full Fact perform manual fact-checking to verify claims in different domains. However, this is a laborious task requiring domain expertise Adair et al. (2017); Hanselowski (2020), which has become increasingly infeasible due to the sheer volume of misinformation generated by humans and AI models. Automated fact-checking has become an increasingly popular approach to verify the veracity of claims in a given text. [^2] We acknowledge that claim verification models can also be applied to other modalities of data (e.g., images). In this survey, we address models applied to text only. There are several steps involved in the fact-verification pipelines; the three main components are claim detection, evidence retrieval, and veracity prediction. Models used in these steps have followed the general methodological developments of the field and we have thus observed an increase in the use of LLMs for claim verification Zhang and Gao (2023); Wang and Shu (2023); Quelle and Bovet (2024). Figure 1 shows an example claim veracity scenario and compares the architectural differences between traditional NLP-based vs LLM-based claim verification systems. The latter are less prone to error propagation, and provide additional justifications when verifying claims.
@@ 1 Introduction
Despite this, LLMs are pre-trained on very large text collections texts and are prone to hallucinations, often generating texts containing incorrect information. Further, such models can be used to generate misinformation at scale Chen and Shu (2023); Zhou et al. (2023); Dmonte et al. (2024) and can therefore be exploited by malicious actors to spread factually incorrect information at an unprecedented rate Pan et al. (2023c). Furthermore, using these pre-trained models for fact-verification may generate incorrect veracity labels, as the models may also rely on obsolete information. Approaches like RAG Gao et al. (2023) are used for the model to retrieve the most recent information during fact verification.
@@
A few general automated claim verification surveys have been published Zeng et al. (2021); Bekoulis et al. (2021); Guo et al. (2022) as well as a couple of surveys on particular aspects of the task such as explainability Vallayil et al. (2023) and applications to specialized domains such as scientific texts Vladika and Matthes (2023). However, all past related surveys lack consideration for LLM-based approaches or focus on specific sub-tasks of the pipeline Panchendrarajan and Zubiaga (2024). In this paper, we fill this important gap by surveying LLM-based frameworks proposed in recent years. To the best of our knowledge, this is the first survey to explore claim verification with LLMs.
2 Search Criteria
@@
We search popular repositories [^3] ACM: portal.acm.org. IEEE Xplore: ieeexplore.ieee.org. Scopus: scopus.com. ACL: aclanthology.org. Web of Science: isiknowledge.com. Springer: link.springer.com. ArXiv: arxiv.org. CEUR: https://ceur-ws.org/. of scientific articles to collect the papers that serve as primary sources for this survey, using the terms LLMs, claim verification, fact verification and related keywords. We focus primarily on the ACL Anthology, ACM Digital Library, IEEE Xplore, and proceedings of related conferences such as AAAI and IJCAI. We further search on Scopus, Springer Link, Science@Direct, and ArXiv.
@@
Based on our search terms, we identify over 100 papers, and filter them as follows: (i) since the scope of our review is text-based LLM claim verification systems, we omit papers related to multimodal approaches (e.g., text-to-images), and papers investigating other modalities (e.g., images, graph-based systems); (ii) since we focus on LLM-based systems for claim veracity, we omit papers on claim identification, claim detection, and/or detecting LLM-generated content. This results in a total of 49 papers related to LLM-based approaches for veracity labeling. The papers on topic LLM-based veracity labeling have been published primarily in the ACL Anthology and ACM Digital Library, but also in other repositories such as the IEEE Xplore.
3 Claim Verification Pipeline
@@
Figure 2 shows a typical claim verification pipeline consisting of: claim detection, claim matching, claim check-worthiness, document/evidence retrieval, rationale/sentence selection, veracity label prediction, and explanation/justification generation. Many systems only make use of some of these modules.
Figure 2: A typical claim verification pipeline
@@
Input texts may contain one or more statements, not all of which are claims. Given the input text, a claim detection module is designed to identify all the statements containing a claim. For example, the statement ’I loved the movie Oppenheimer.’ is an opinion, whereas the statement ’The COVID-19 pandemic started in Texas.’ contains a claim.
@@
Not all identified claims are check-worthy. Given an input claim, the task is to identify the claims that include real-world assertions and that may need to be verified Hassan et al. (2015); Nakov et al. (2021). Check-worthy claim identification is a subjective task and it relies on factors like popularity of the claim, public interest to determine the claim’s veracity, etc. For example, the claim ’The President met the State Governor to discuss the infrastructure deal.’ is less check-worthy than the claim ’Drinking salt water cures COVID.’
@@
The identified check-worthy claims can be matched to previously fact-checked claims. Given an input claim and a database of previously fact-checked claims, claim matching is used to determine if the input claim is previously fact-checked and exists in the database Shaar et al. (2020); Nakov et al. (2021). This can help predict the veracity label directly without additional steps.
@@
If the input claim does not exist in the database of fact-checked claims, it needs to be verified. In the evidence retrieval sub-task, all relevant documents related to the input claim are extracted, either from an external database or through an internet search Chen et al. (2017). The threshold of the number of documents to be retrieved can be predetermined.
@@
Not all information in the retrieved documents is relevant to the claim. Hence for the rationale selection task, only the information or evidence most relevant to the claim is selected to predict the veracity label.
@@
Once a rationale is selected, it is provided to a classifier along with the claim and potentially additionally features, to predict a veracity label; often ‘SUPPORTED’, ‘REFUTED’, vs ‘NOT ENOUGH EVIDENCE’. For some datasets, the labels can be ‘TRUE’ vs ‘FALSE’.
@@
Recent works have focused on generating explanations for the veracity labels prediction. This specific task is focused on generating natural language justifications or explanations for the prediction, considering the claim and evidence.
4 LLM Approaches
@@
Figure 3 shows an example pipeline that encapsulates multiple component modules (i.e., Evidence Retrieval, Prompt Creation, Transfer Learning, and LLM Generation) for verifying claims. Different from traditional fact verification pipelines that select evidence sets for verifying claims, LLM-based claim verification conditions generated text based on the concatenated input claim and retrieved evidence. Retrieval-augmented LLMs have been shown to perform well on knowledge-intensive tasks such as text generation.
Figure 3: LLM-based claim verification pipeline. This involves creating a prompt from the retrieved evidence and the input claim as input to the LLM to generate a label, sentence evidence, and/or explanation of its response.
4.1 Evidence Retrieval Strategies
@@
RAG models, which have been developed to address hallucination in LLMs for knowledge-intensive tasks, have shown success for fact verification Gao et al. (2023); Guan et al. (2023). Early work such as Lewis et al. (2020) develop a framework that retrieves evidence from an external database such as Wikipedia for conditionally generating veracity labels. Izacard et al., Izacard et al. (2023) demonstrate that RAGs perform well on the FEVER shared task Thorne et al. (2018b)) in few-shot settings, showing 5% improvement over large-scale LLMs such as Gopher Rae et al. (2021) with significantly fewer parameters (i.e., 11B compared to 280B). Other works consider the optimization of either document ranking Glass et al. (2022); Chen et al. (2022c) or input claims (i.e., queries) Hang et al. (2024) as a crucial step for improving evidence retrieval for veracity labeling. Hofstätter et al., Hofstätter et al. (2023) use an autoregressive re-ranker to get the most relevant passages from the retriever. These are then passed to the generation model to generate the veracity label.
@@
Despite this, RAG models often fail when encountering long or complex input claims, resulting in incorrectly generated veracity labels or evidence sentences. Recent works have addressed this by segmenting long claims into smaller sub-claims and performing multiple rounds of retrieval. Khattab et al., Khattab et al. (2021) presents a pipeline for multi-hop claim verification that uses an iterative retriever and neural methods for effective document retrieval and re-ranking. Shao et al. Shao et al. (2023) show that using a re-ranker to distill knowledge to a retriever helps close the semantic gaps between a query and document passage when verifying claims using iterative RAGs. Other works address the issue of complex claims using fine-grained retrieval techniques based on claim decomposition. Ches et al. Chen et al. (2023b) generate sub-questions based on a claim, which a document retriever uses to retrieve relevant documents. A fine-grained retriever retrieves top-k text spans as evidence based on a k-word window and BM25. Zhang and Gao Zhang and Gao (2023) decompose a claim into sub-claims and generate questions to verify the sub-claims. External knowledge sources are used to retrieve relevant information to verify the sub-claims and generate a final veracity label. Pan et al. Pan et al. (2023b) follow a programming paradigm, where the claim is broken down into subtasks and the final label is the aggregation of the execution of each subtask. The fact-verification subtask uses external knowledge to retrieve relevant evidence for a claim.
@@
Hang et al. Hang et al. (2024) retrieve evidence based on generated knowledge graphs of evidences. They generate a knowledge graph of the user query or input claim and compare it to the database of knowledge graphs to retrieve the most relevant information for claim verification. Hu et al. Hu et al. (2023) propose a latent variable model that allows the retrieval of the most relevant evidence sentences from a document while removing the irrelevant sentences. This approach reduces noisy data during the verification process. Stochastic-RAG proposed by Zamani and Bendersky Zamani and Bendersky (2024) uses a stochastic sampling without replacement process for evidence retrieval and selection. This approach overcomes the ranking and selection of the evidence hence optimizing the RAG model. Zhang et al. Zhang et al. (2023) optimize the evidence retrieval process by using feedback from the claim verifier. The divergence between the evidence from a retrieved evidence set provided to the verifier and the gold standard evidence, acts as a feedback signal used to train the retriever. Xu et al. Xu et al. (2024b) propose Search-in-the-Chain, where an LLM generates a reasoning chain and based on the answer to each node in the chain, the retrieval can be used to correct an answer or provide additional knowledge, improving the generation accuracy of an LLM.
4.2 Prompt Creation Strategies
@@
Text prompting has shown to be effective for improving the output of LLMs. In the context of claim verification, several works investigate both manual and automated prompting strategies to increase robustness. Zhang and Gao Zhang and Gao (2023) develop a hierarchical prompting technique to verify multiple sub-claims using a step-by-step approach. Li et al. Li et al. (2023c) demonstrates a self-sufficient claim verification through prompting instructions on multiple language models.
@@
ProToCo Zeng and Gao (2023) demonstrates improved claim verification performance by leveraging a consistency mechanism to construct variants of the original claim-evidence pair prompt based on three logical relations (i.e., confirmation, negation, uncertainty). Chen et al. Chen et al. (2023a) develop a unified retrieval framework that employs discrete, continuous, and hybrid prompt strategies for adjusting to various knowledge-intensive tasks such as claim verification. The FactualityPrompts Lee et al. (2022) framework tests the output generations of LLMs given an input prompt and uses an external database such as Wikipedia to calculate factuality and quality measures compared to the ground truth. Other works aim to improve an LLM’s reasoning abilities by appending the claims with evidence during prompting for claim verification and text generation Parvez (2024); Dougrez-Lewis et al. (2024).
4.3 Transfer Learning Strategies